Use the Rational Zeros Theorem to write a list of all possible rational zeros of the function.

f(x) = -2x4 + 4x3 + 3x2 + 18


1
Expert's answer
2021-09-22T23:55:53-0400

If the rational number "\frac{a}{b}," where "a\in\Z" and "b\in\N," is a rational zero of the polinomial "f(x) = -2x^4 + 4x^3 + 3x^2 + 18," then according to the Rational Zeros Theorem "a" is a divisor of 18 and "b" is a divisor of "-2." Then all possible zeros are the following:

"1,\ -1,\ 2,\ -2,\ 3,\ -3,\ 6,\ -6,\ 9,\ -9,\ 18,\ -18,\ \frac{1}2,\ -\frac{1}2,\ \frac{3}2,\ -\frac{3}2,\ \frac{9}2,\ -\frac{9}2."